just xiii thanks
(xi) How many different 5-card poker hands can be dealt from a regular 52-card (xii) How many of these hands contain no aces? (xii) How many contain a aces, for a-0 to 4? (xiv) How many contain all cards of the same suit? deck?
Answer #1

xi)

Possible Poker Hands in 52-Card Deck. The total number of possible hands can be found by taking the combination between 52 and 5. That is the number of ways 5 cards can be chosen out of 52 cards.  This means that if there are 52 cards, how many combinations of 5 cards can be drawn in 2 3ways.

That is,

(52)-而21丽 471 x ริเ = 2598960 2598960

xii)

if there is no Ace, that means the cards have to be dealt from a deck where there is no ace. he total number of possible hands can be found by taking the combination between 48 and 5. That is the number of ways 5 cards can be chosen out of 48 cards.  This means that if there are 52 cards, how many combinations of 5 cards can be drawn in 45ways.

That is,

( ) 48 = 4748,5,- 1712304 ,-, ー

xiii)

The required answer is,

/52 times6+/52 /51 /49

xiv)

Five cards of the same suit mean the cards will be from either of the four suits. The required number of ways is,

(13 ) 13 ×4=1287